# Lagrangian extensions and left symmetric structures on the four-dimensional real Lie superalgebras

### v1

@@ 2.3. The 4-dimensional real Lie superalgebras
    Table 1 having the trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=2|2$;
@@
    Table 2 having the non-trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=2|2$;
@@
    Table 3 having the trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=3|1$;
@@
    Table 4 having the trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=1|3$;
@@
    Table 5 having the non-trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=3|1$;
@@
    Table 6 having the non-trivial algebras, i.e. $[\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=1|3$.
In this paper, we present only the forms as in Table 1-6 which are suitable for Lagrangian extensions, as opposed to [BM], where the forms are given in the form of parametric families. The classification of these structures up to a symplectomorphism is still an open problem. We also correct a claim in [BM] stating that only non-homogenous forms exist for the Lie superalgebras $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$. It turns out that these two superalgebras both admit an even and an odd non-degenerate closed form, and both of them can be obtained as Lagrangian extensions.
3. The notion of $T^{*}$-extensions and $\Pi T^{*}$- (Lagrangian extensions)
We will introduce the notion of $T^{*}$-extension and $\Pi T^{*}$-extension of Lie superalgebras (Lagrangian extensions). Originally, the notion of $T^{*}$-extension for Lie algebras was first introduced by Bordemann in [Bor]. Recently, it was studied in [BC] in the context of quasi-Frobenius Lie algebras, and then in [BM] in the context of quasi-Frobenius Lie superalgebras.
@@ 3.1. Flat connections and representations
Let $({\mathfrak{h}},\nabla)$ be a Lie superalgebra endowed with a flat connection $\nabla$. Since the connection $\nabla$ is flat, the map
@@
defines a representation. We define a map $\rho:\mathfrak{h}\rightarrow\mathrm{End}(\mathfrak{h}^{*})$, the dual representation, as follows (for all $u\in\mathfrak{h}$ and $\xi\in\mathfrak{h}^{*}$):
@@ 3.2. Polarization and Lagrangian extensions
Following [BC] (for the supercase, see [BM]), a polarization for a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is a choice of a Lagrangian subalgebra $\mathfrak{l}$ of $(\mathfrak{g},\omega)$. A strong polarization of a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is a pair $(\mathfrak{a},N)$ consisting of a Lagrangian ideal $\mathfrak{a}\subset\mathfrak{g}$ and a complementary Lagrangian subspace $N\subset\mathfrak{g}$. The quadruple $(\mathfrak{g},\omega,\mathfrak{a},N)$ is referred to as a strongly polarized quasi-Frobenius Lie superalgebra.
@@
Recall that the space $\mathfrak{h}^{*}$ and $\Pi(\mathfrak{h}^{*})$ are $\mathfrak{h}$-modules by means of the representations (6) and (7).
@@
An even form $\omega$ is defined on $\mathfrak{g}$ as follows: (for all $u+\xi,v+\zeta\in\mathfrak{h}\oplus\mathfrak{h}^{*}$)
(8) $\omega(u+\xi,v+\zeta)=\xi(v)-(-1)^{|\zeta||u|}\zeta(u).$
@@
We define an odd form $\kappa$ on $\mathfrak{k}$ as follows: (for all $u+\Pi(\xi),v+\Pi(\zeta)\in\mathfrak{h}\oplus\Pi(\mathfrak{h}^{*}$))
(9) $\kappa(u+\Pi(\xi),v+\Pi(\zeta))=\xi(v)-(-1)^{(|\zeta|+1)|u|}\zeta(u).$
@@
(10) $(-1)^{|u||w|}\alpha(u,v)(w)+\circlearrowleft(u,v,w)=0\text{ for all $u,v,w\in\mathfrak{h}$}.$
@@
(11) $(-1)^{|u||w|}\Pi\circ\beta(u,v)(w)+\circlearrowleft(u,v,w)=0\text{ for all $u,v,w\in\mathfrak{h}$}.$
@@
We arrive to the following theorem proved in [BM].
@@
The converse of this theorem says that every orthosymplectic or periplectic quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ that has a Lagrangian ideal $\mathfrak{a}$ can be obtained as a $T^{*}$-extension or $\Pi T^{*}$-extension of a flat Lie superalgebra.
Let $(\mathfrak{g},\omega,\mathfrak{a},N)$ be a strongly polarized quasi-Frobenius Lie superalgebra. Here the form $\omega$ is either orthosymplectic or periplectic. Consider the quotient space $\mathfrak{h}=\mathfrak{g}/\mathfrak{a}$ (recall that $\mathfrak{a}^{\perp}=\mathfrak{a}$). We will construct a flat connection $\nabla$ on $\mathfrak{h}$ (see [BM]), superizing the construction of [BC]. For every $u,v\in\mathfrak{h}$, we denote by $\tilde{u}$ and $\tilde{v}$ their lift to $\mathfrak{g}$. We then write
@@ 4. The 4-dimensional Lie superalgebras as Lagrangian extensions
    The Lie superalgebra $D^{5}$: We correct here the description of the Lie superalgebra $D_{5}$ in [BM]. It is $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=e_{4}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ on $\mathfrak{h}$ is given by:
@@
    $00=\omega(v,e_{4})=-l_{2}\implies l_{2}=$ 0 .
    This implies that $\mathfrak{a}^{\bot}=\mathfrak{a}$. The Lie superalgebra $\mathfrak{h}$ is given by $\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $[e_{1},e_{2}]_{\mathfrak{h}}=0$. The connection $\nabla$ is given by:
@@
    Now, we compute the torsion and curvature of $\nabla$, using Eqns. (4, 5). For the torsion, we have:
    $T(e_{1},e_{2}):=\nabla_{e_{1}}(e_{2})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}(e_{1})-[e_{1},e_{2}]_{\mathfrak{h}}=-e_{2}-(-e_{2})-0=0.$
@@
    $\begin{array}[]{lcl}R^{\nabla}(e_{1},e_{2})(e_{1})&:=&\nabla_{e_{1}}\nabla_{e_{2}}(e_{1})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}\nabla_{e_{1}}(e_{1})-\nabla_{[e_{1},e_{2}]_{\mathfrak{h}}}(e_{1})\\[5.69054pt] &=&\nabla_{e_{1}}(-e_{2})-\nabla_{e_{2}}(-e_{1})=e_{2}-e_{2}=0.\end{array}$
@@
    $\begin{array}[]{lcl}R^{\nabla}(e_{1},e_{2})(e_{2})&:=&\nabla_{e_{1}}\nabla_{e_{2}}(e_{2})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}\nabla_{e_{1}}(e_{2})-\nabla_{[e_{1},e_{2}]_{\mathfrak{h}}}(e_{2})\\[5.69054pt] &=&\nabla_{e_{1}}(e_{1})-\nabla_{e_{2}}(-e_{2})=-e_{1}+e_{1}=0.\end{array}$
@@
    The Lie superalgebra $D^{7}_{-1q}$, with $q\leq 1$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $[e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{4}=\Pi(e_{1}^{*}),e_{3}=\Pi(e_{2}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
@@
    The Lie superalgebra $D^{7}_{pq}$, with $p=-q,q\neq 0,-1$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=-pe_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
@@
    The Lie superalgebra $D^{7}_{pq}$, with $p=-q,q=-1$: In the case where the form is odd, then this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $[e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{3}=\Pi(e_{1}^{*}),e_{4}=\Pi(e_{2}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
@@
    In the case where the form is even, then this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=-e_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
@@
    The Lie superalgebra $D^{8}_{-1}$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $[e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{3}=\Pi(e_{1}^{*}),e_{4}=\Pi(e_{2}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
@@
    The Lie superalgebra $D^{10}_{q}$ with $q\not=-1$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=qe_{4}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{2}=\Pi(e_{4}^{*})|e_{3}=-(q+1)\Pi(e_{1}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by:
@@
[…]
@@
      If $x_{3}=0,$ and $x_{4}\neq 0$. Looking at brackets again we get $[e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{4}e_{4},$ and $[e_{4},X]=x_{4}e_{2}$, which implies $e_{2},e_{4}\in\mathfrak{a}$, so $\mathfrak{a}=\mathrm{Span}\{e_{2},e_{4}\}$. We have to check that $\mathfrak{a}$ is not Lagrangian. Indeed, $\omega(e_{4},e_{4})=1$.
@@
      If $x_{3}\neq 0,$ and $x_{4}\neq 0$. Since all the previous cases showed that we can’t find a Lagrangian ideal, and since the proof is symmetrical in X and Y without loss of generality we can state that $y_{3}\neq 0,y_{4}\neq 0$. This implies that $e_{2}\in\mathfrak{a}$ and that $x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3}+\dfrac{1}{2}x_{4}e_{4}\in\mathfrak{a}$. We can write $X=e_{2}$ and $Y=z_{3}e_{3}+z_{4}e_{4}$. Once again, $\omega(Y,Y)=z_{3}^{2}+z_{4}^{2}\not=0$ which implies that $\mathfrak{a}$ is not Lagrangian.
@@
[…]
@@
    The Lie superalgebra $(D^{7}_{1-p,p})$, with $p\leq\dfrac{1}{2}$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $[e_{1},e_{4}]_{\mathfrak{h}}=(1-p)e_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by:
@@
    The Lie superalgebra $(D^{9}_{1/2,p})$, with $p>0$: This Lie superalgebra does not admit a 2-dimensional Lagrangian ideal. Indeed, let us suppose there exists one, denoted by $\mathfrak{a}=\mathrm{Span}\{X,Y\}$. Let us write $X=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}+x_{4}e_{4}$ and $Y=y_{1}e_{1}+y_{2}e_{2}+y_{3}e_{3}+y_{4}e_{4}$. We compute the brackets $[e_{i},X]$ for all $i\in\{1,2,3,4\}$. We get
@@
    We will analyze a few cases. If $x_{1}\neq 0$, given the second bracket we get that $e_{2}\in\mathfrak{a}$ and according to the last brackets $e_{3}\in\mathfrak{a}$ and $e_{4}\in\mathfrak{a}$. This case is not possible since $\mathfrak{a}$ has dimension 2. This implies that $x_{1}=0$. We get then:

### v2

@@ 2.3. The 4-dimensional real Lie superalgebras
    Table 1 having the trivial algebras, i.e. $\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=2|2$;
@@
    Table 2 having the non-trivial algebras, i.e. $\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=2|2$;
@@
    Table 3 having the trivial algebras, i.e. $\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=3|1$;
@@
    Table 4 having the trivial algebras, i.e. $\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]=\{0\}$ and with $\mathrm{sdim}=1|3$;
@@
    Table 5 having the non-trivial algebras, i.e. $\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=3|1$;
@@
    Table 6 having the non-trivial algebras, i.e. $\mathfrak{g_{\bar{1}}},\mathfrak{g_{\bar{1}}}]\not=\{0\}$ and with $\mathrm{sdim}=1|3$.
In this paper, we present only the forms as in Table 1-6 which are suitable for Lagrangian extensions, as opposed to [BM], where the forms are given in the form of parametric families. The classification of these structures up to a symplectomorphism is still an open problem. We also correct a claim in [BM] stating that only non-homogeneous forms exist for the Lie superalgebras $(D_{0}^{10})^{1}$ and $(D_{0}^{10})^{2}$. It turns out that these two superalgebras both admit an even and an odd non-degenerate closed form, and both of them can be obtained as Lagrangian extensions. We only consider indecomposable Lie superalgebras.
3. The notion of $T^{*}$- extensions and $\Pi T^{*}$- extensions
We will introduce the notion of $T^{*}$-extension and $\Pi T^{*}$-extension of Lie superalgebras (Lagrangian extensions). The notion of a $T^{*}$-extension for Lie algebras was originally introduced by Bordemann in [Bor]. Recently, it was studied in [BC] in the context of quasi-Frobenius Lie algebras, and then in [BBE, BM] in the context of quasi-Frobenius Lie superalgebras.
@@ 3.1. Flat connections and representations
Let $({\mathfrak{h}},\nabla)$ be a Lie superalgebra endowed with a flat connection $\nabla$. Since the connection $\nabla$ is flat, it follows that the map
@@
defines a representation. We define a map $\rho:\mathfrak{h}\rightarrow\mathrm{End}(\mathfrak{h}^{*})$, the dual representation, as follows: (for all homogeneous $u\in\mathfrak{h}$ and $\xi\in\mathfrak{h}^{*}$)
@@
These two representations will be used to build Lagrangian extensions in the next subsection.
@@ 3.2. Polarization and Lagrangian extensions
Following [BC] (for the supercase, see [BBE, BM]), a polarization for a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is a choice of a homogeneous Lagrangian subalgebra $\mathfrak{l}$ of $(\mathfrak{g},\omega)$ (namely, $\mathfrak{l}=\mathfrak{l}^{\perp}$, with $\mathfrak{l}^{\perp}$ the orthogonal with respect to $\omega$). A strong polarization of a quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ is a pair $(\mathfrak{a},N)$ consisting of a homogeneous Lagrangian ideal $\mathfrak{a}\subset\mathfrak{g}$ and a complementary Lagrangian subspace $N\subset\mathfrak{g}$. Following [BC], the quadruple $(\mathfrak{g},\omega,\mathfrak{a},N)$ is referred to as a strongly polarized quasi-Frobenius Lie superalgebra.
The following construction is borrowed from [BM], as a generalization to [BC].
@@
Recall that the two spaces $\mathfrak{h}^{*}$ and $\Pi(\mathfrak{h}^{*})$ are $\mathfrak{h}$-modules by means of the representations (3) and (4).
@@
An even form $\omega$ is defined on $\mathfrak{g}$ as follows
(5) $\omega(u+\xi,v+\zeta):=\xi(v)-(-1)^{|\zeta||u|}\zeta(u),$
for all homogeneous $u+\xi,v+\zeta\in\mathfrak{h}\oplus\mathfrak{h}^{*}$.
@@
We define an odd form $\kappa$ on $\mathfrak{k}$ as follows
(6) $\kappa(u+\Pi(\xi),v+\Pi(\zeta)):=\xi(v)-(-1)^{(|\zeta|+1)|u|}\zeta(u),$
for all homogeneous $u+\Pi(\xi),v+\Pi(\zeta)\in\mathfrak{h}\oplus\Pi(\mathfrak{h}^{*})$.
@@
(7) $(-1)^{|u||w|}\alpha(u,v)(w)+\circlearrowleft(u,v,w)=0\text{ for all homogeneous $u,v,w\in\mathfrak{h}$},$
where here and in what follows, the symbol $\circlearrowleft(u,v,w)$ denotes the sum over cyclic permutations of the variables $u,v,w$.
@@
(8) $(-1)^{|u||w|}\Pi\circ\beta(u,v)(w)+\circlearrowleft(u,v,w)=0\text{ for all homogeneous $u,v,w\in\mathfrak{h}$}.$
@@
We arrive to the following theorem proved in [BM] (see also [BC] in the non-super case).
@@
The converse of this theorem says that every orthosymplectic or periplectic quasi-Frobenius Lie superalgebra $(\mathfrak{g},\omega)$ that has a homogeneous Lagrangian ideal $\mathfrak{a}$ can be obtained as a $T^{*}$-extension or $\Pi T^{*}$-extension of a flat Lie superalgebra.
Let $(\mathfrak{g},\omega,\mathfrak{a},N)$ be a strongly polarized quasi-Frobenius Lie superalgebra. Here the form $\omega$ is either orthosymplectic or periplectic. Consider the quotient space $\mathfrak{h}=\mathfrak{g}/\mathfrak{a}$ (recall that $\mathfrak{a}^{\perp}=\mathfrak{a}$). We will construct a flat connection $\nabla$ on $\mathfrak{h}$ (see [BM]), superizing the construction of [BC]. For every homogeneous $u,v\in\mathfrak{h}$, we denote by $\tilde{u}$ and $\tilde{v}$ their lift to $\mathfrak{g}$. We then write
@@
The following theorem is needed to determine whether the Lie superalgebras in Backhouse’s list arise as Lagrangian extensions of smaller Lie superalgebras. We refer to [BM] for its proof.
@@ 4. The 4-dimensional Lie superalgebras as Lagrangian extensions
    The Lie superalgebra $D^{5}$: We correct here the description of the Lie superalgebra $D_{5}$ in [BM]. It is $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{1},e_{4}]_{\mathfrak{h}}=e_{4}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{2}=-\Pi(e_{4}^{*})|e_{3}=\Pi(e_{1}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ on $\mathfrak{h}$ is given by:
@@
    $0=\omega(v,e_{4})=-l_{2}\implies l_{2}=0.$
    This implies that $\mathfrak{a}^{\bot}=\mathfrak{a}$. The Lie superalgebra $\mathfrak{h}$ is given by $\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $e_{1},e_{2}]_{\mathfrak{h}}=0$. The connection $\nabla$ is given by
@@
    Now, we compute the torsion and curvature of $\nabla$, using Eqns. (1, 2). For the torsion, we have
    $T(e_{1},e_{2})=\nabla_{e_{1}}(e_{2})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}(e_{1})-[e_{1},e_{2}]_{\mathfrak{h}}=-e_{2}-(-e_{2})-0=0.$
@@
    $\begin{array}[]{lcl}R^{\nabla}(e_{1},e_{2})(e_{1})&=&\nabla_{e_{1}}\nabla_{e_{2}}(e_{1})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}\nabla_{e_{1}}(e_{1})-\nabla_{[e_{1},e_{2}]_{\mathfrak{h}}}(e_{1})\\[5.69054pt] &=&\nabla_{e_{1}}(-e_{2})-\nabla_{e_{2}}(-e_{1})=e_{2}-e_{2}=0.\end{array}$
@@
    $\begin{array}[]{lcl}R^{\nabla}(e_{1},e_{2})(e_{2})&=&\nabla_{e_{1}}\nabla_{e_{2}}(e_{2})-(-1)^{|e_{1}||e_{2}|}\nabla_{e_{2}}\nabla_{e_{1}}(e_{2})-\nabla_{[e_{1},e_{2}]_{\mathfrak{h}}}(e_{2})\\[5.69054pt] &=&\nabla_{e_{1}}(e_{1})-\nabla_{e_{2}}(-e_{2})=-e_{1}+e_{1}=0.\end{array}$
@@
    The Lie superalgebra $D^{7}_{-1q}$, with $q\leq 1$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{4}=\Pi(e_{1}^{*}),e_{3}=\Pi(e_{2}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by
@@
    The Lie superalgebra $D^{7}_{pq}$, with $p=-q,q\neq 0,-1$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{1},e_{4}]_{\mathfrak{h}}=-pe_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
@@
    The Lie superalgebra $D^{7}_{pq}$, with p=q,q=1p=-q,q=-1: In the case where the form is odd, then this Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{3}=\Pi(e_{1}^{*}),e_{4}=\Pi(e_{2}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by
@@
    In the case where the form is even, then this Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{1},e_{4}]_{\mathfrak{h}}=-e_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
@@
    The Lie superalgebra $D^{8}_{-1}$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1},e_{2}\}$ with the bracket $e_{1},e_{2}]_{\mathfrak{h}}=e_{2}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{3}=\Pi(e_{1}^{*}),e_{4}=\Pi(e_{2}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by
@@
    The Lie superalgebra $D^{10}_{q}$ with $q\not=-1$: This Lie superalgebra is a $\Pi T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{1},e_{4}]_{\mathfrak{h}}=qe_{4}$, and where $\Pi\mathfrak{h}^{*}$ is spanned by $e_{2}=\Pi(e_{4}^{*})|e_{3}=-(q+1)\Pi(e_{1}^{*})$. The 2-form $\beta\equiv 0$ and the connection $\nabla$ is given by
@@
    The Lie superalgebra $(D^{7}_{1/2\;1/2})^{1}$: This Lie superalgebra does not admit a 2-dimensional Lagrangian ideal. Indeed, let us suppose the contrary, and let us denote by $\mathfrak{a}=\mathrm{Span}\{X,Y\}$ such an ideal. Let us write $X=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}+x_{4}e_{4}$ and $Y=y_{1}e_{1}+y_{2}e_{2}+y_{3}e_{3}+y_{4}e_{4}$. We compute the brackets $e_{i},X]$, for all $i\in\{1,2,3,4\}$. we obtain
@@
      If $x_{3}=0,$ and $x_{4}\neq 0$. Looking at brackets again we get $e_{1},X]=x_{2}e_{2}+\dfrac{1}{2}x_{4}e_{4},$ and $e_{4},X]=x_{4}e_{2}$, which implies $e_{2},e_{4}\in\mathfrak{a}$, so $\mathfrak{a}=\mathrm{Span}\{e_{2},e_{4}\}$. We have to check that $\mathfrak{a}$ is not Lagrangian. Indeed, $\omega(e_{4},e_{4})=1$.
@@
      If $x_{3}\neq 0,$ and $x_{4}\neq 0$. Since all the previous cases showed that we cannot find a Lagrangian ideal, and since the proof is symmetrical in X and Y without loss of generality we can state that $y_{3}\neq 0,y_{4}\neq 0$. This implies that $e_{2}\in\mathfrak{a}$ and that $x_{2}e_{2}+\dfrac{1}{2}x_{3}e_{3}+\dfrac{1}{2}x_{4}e_{4}\in\mathfrak{a}$. We can write $X=e_{2}$ and $Y=z_{3}e_{3}+z_{4}e_{4}$. Once again, $\omega(Y,Y)=z_{3}^{2}+z_{4}^{2}\not=0$ which implies that $\mathfrak{a}$ is not Lagrangian.
@@
    The Lie superalgebra $(D^{7}_{1/2\;1/2})^{2}$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|x:=e_{3}-e_{4}\}$ with the bracket $e_{1},x]_{\mathfrak{h}}=\dfrac{1}{2}x$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=e_{1}^{*}|e_{3}+e_{4}=-2x^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
@@
    The Lie superalgebra $(D^{7}_{1-p,p})$, with $p\leq\dfrac{1}{2}$: This Lie superalgebra is a $T^{*}$-extension of the Lie superalgebra $\mathfrak{h}=\mathrm{Span}\{e_{1}|e_{4}\}$ with the bracket $e_{1},e_{4}]_{\mathfrak{h}}=(1-p)e_{4}$, and where $\mathfrak{h}^{*}$ is spanned by $e_{2}=-e_{1}^{*}|e_{3}=e_{4}^{*}$. The 2-form $\alpha\equiv 0$ and the connection $\nabla$ is given by
@@
    The Lie superalgebra $(D^{9}_{1/2,p})$, with $p>0$: This Lie superalgebra does not admit a 2-dimensional Lagrangian ideal. Indeed, let us suppose there exists one, denoted by $\mathfrak{a}=\mathrm{Span}\{X,Y\}$. Let us write $X=x_{1}e_{1}+x_{2}e_{2}+x_{3}e_{3}+x_{4}e_{4}$ and $Y=y_{1}e_{1}+y_{2}e_{2}+y_{3}e_{3}+y_{4}e_{4}$. We compute the brackets $e_{i},X]$, for all $i\in\{1,2,3,4\}$. We obtain
@@
    We will analyze a few cases. If $x_{1}\neq 0$, given the second bracket we get that $e_{2}\in\mathfrak{a}$ and according to the last brackets $e_{3}\in\mathfrak{a}$ and $e_{4}\in\mathfrak{a}$. This case is not possible since $\mathfrak{a}$ has dimension 2. This implies that $x_{1}=0$. We get then